How many triangles are possible with I gikai angle A=1350, a 12m, b = 15 m²
Answer #1

2. 2 = A = 135 a = 12m b=15m. a С Sin SinB Soc 12 15 С Süs (135) SinB sist Sis B = 158 0.7071 = 0.8838 12 B= sin(0-8838 07 Bo C= 180- 197.1 c =- 17.10 This is not possible because sum of all angles of a trangle should be 180So, no triangle is possib